COLLISIONAL QUENCHING OF OH A^{2}Σ^{+} BY H_{2} AND N_{2}: DYNAMICAL OUTCOMES

LOGAN P. DEMPSEY, TIMOTHY D. SECHLER, CRAIG MURRAY, \underline{MARSHA I. LESTER}, Department of Chemistry, University of Pennsylvania, Philadelphia, PA 19104-6323; SPIRI-DOULA MATSIKA, Department of Chemistry, Temple University, Philadelphia, Pennsylvania 19122-6014.

The nascent OH X^{2}Π product state distributions arising from collisional quenching of electronically excited OH A^{2}Σ^{+} by H_{2} and N_{2} have been determined using a pump-probe technique. For both collision partners, the majority of OH X^{2}Π products are observed in their lowest vibrational level, v′′ = 0, with significantly less population in v′′ = 1. The OH (v′′ = 0) products are generated with a substantial degree of rotational excitation, peaking around N′′ = 15 with H_{2} as the collision partner and N′′ = 18 with N_{2}. Complementary measurements of the branching fraction into OH X^{2}Π product states demonstrate that reaction is the dominant decay pathway for quenching of OH A^{2}Σ^{+} by H_{2}, while nonreactive quenching is the dominant pathway for N_{2}. These observations are discussed in the context of theoretical calculations that examine the topography of the conical intersections which couple the electronically excited and ground state potential energy surfaces. The experimental observables are interpreted as dynamical signatures of nonadiabatic passage through the conical intersection regions responsible for quenching in both systems.